OpenStudy (anonymous):

Evaluate tan(sec^-1 X). How do i simplify the secant using a triangle?

OpenStudy (anonymous):

\[\Huge \sec(\theta)=\frac{ x }{ 1 }=x\]\[\Huge \sec^{-1}(x)=\theta\]Thus\[\Huge \tan(\sec^{-1}(x))=\tan(\theta)=\frac{ \sqrt{x^{2}-1} }{ 1 }\]

OpenStudy (anonymous):

Thank You ^^